) Given, E ̅ = E ̅m sin(x) sin (t) a ̂y and H ̅ = E ̅m/µo cos(x) cos (t) a ̂z. Determine whether these fields satisfy Maxwell’s equations or not.
1
Expert's answer
2020-09-15T10:04:14-0400

Let's substitute them to the Maxwell’s equations (with no charges and currents in vacuum):


"\nabla\cdot\mathbf{D}= 0\\
\nabla\cdot\mathbf{B}=0\\
\nabla\times\mathbf{E}=-\frac{\partial\mathbf{B}}{\partial t}\\
\nabla\times\mathbf{H}= \frac{\partial\mathbf{D}}{\partial t}"

Considering the fact that "\mathbf{D} = \varepsilon_0\mathbf{E}" and "\mathbf{B} = \mu_0\mathbf{H}", obtain:

1.

"\nabla\cdot\mathbf{D}= \varepsilon_0 \left ( \dfrac{\partial E_x}{\partial x} +
\dfrac{\partial E_y}{\partial y}+
\dfrac{\partial E_z}{\partial z} \right)"

As far as "\mathbf{E} = (0,E_m\sin(x)sin(t),0)", y-coordinate does not depend on y, so the divergence is equal to 0: "\nabla\cdot\mathbf{D}=0". The first equation is satisfied.


2. Similarlly, as far as "\mathbf{H} = (0,0,E_m/\mu_0 \cos(x), \cos(t))", the z-coordinate does not depend on z, thus, the divergens will be equal to 0: "\nabla\cdot\mathbf{B} = \nabla\cdot\mathbf{H}=0". The second equation is satisfied.


3. The curle is:


"\nabla\times\mathbf{E}= \left(\frac{\partial E_z}{\partial y} - \frac{\partial E_y}{\partial z}\right) \mathbf e_x+
\left(\frac{\partial E_x}{\partial z} - \frac{\partial E_z}{\partial x}\right) \mathbf e_y+
\left(\frac{\partial E_y}{\partial x} - \frac{\partial E_x}{\partial y}\right) \mathbf e_z"

From all these derivatives only "\dfrac{\partial E_y}{\partial x}" will not be equal to 0. Let's calculate it:


"\dfrac{\partial E_y}{\partial x} =E_m\sin(t)\dfrac{\partial }{\partial x}\sin(x) = E_m\sin(t)\cos(x)"

Thus:


"\nabla\times\mathbf{E}= (0,0,E_m\sin(t)\cos(x) )"

On the other hand:


"-\dfrac{\partial\mathbf{B}}{\partial t} =-\mu_0\dfrac{\partial\mathbf{H}}{\partial t}= 
\left(0,0, - \dfrac{\partial}{\partial t} [E_m \cos(x), \cos(t)] \right)=\\
=(0,0, E_m\sin(t)\cos(x))"

Thus, "\nabla\times\mathbf{E}=-\frac{\partial\mathbf{B}}{\partial t}" and the third equation is also satisfied.


4. The curle is:


"\nabla\times\mathbf{H}= \left(\frac{\partial H_z}{\partial y} - \frac{\partial H_y}{\partial z}\right) \mathbf e_x+
\left(\frac{\partial H_x}{\partial z} - \frac{\partial H_z}{\partial x}\right) \mathbf e_y+
\left(\frac{\partial H_y}{\partial x} - \frac{\partial H_x}{\partial y}\right) \mathbf e_z"

From all these derivatives only "-\dfrac{\partial H_z}{\partial x}" will not be equal to 0. Let's calculate it:


"-\dfrac{\partial H_z}{\partial x} =-E_m/\mu_0\cos(t)\dfrac{\partial }{\partial x}\cos(x) =E_m/\mu_0\cos(t)\sin(x)"

Thus:


"\nabla\times\mathbf{H}= \dfrac{1}{\mu_0} (0,E_m\cos(t)\sin(x),0)"

On the other hand:


"\dfrac{\partial\mathbf{D}}{\partial t} =\varepsilon_0\dfrac{\partial\mathbf{E}}{\partial t}= 
\varepsilon_0\left(0, \dfrac{\partial}{\partial t} [E_m \sin(x), \sin(t)],0 \right)=\\
=\varepsilon_0(0,E_m\cos(t)\sin(x),0)"

Because of the coeffitients "\dfrac{1}{\mu_0}" and "\varepsilon_0" the right hand side is not equal to the left hand side:


"\nabla\times\mathbf{H}\ne \frac{\partial\mathbf{D}}{\partial t}"

Thus, the fourth equation is NOT satisfied.


Answer. These fields do not satisfy Maxwell’s equations.